# GroupNL: Low-Resource and Robust CNN Design over Cloud and Device

### v1

@@ 3.2 Preliminaries of Nonlinear Transformation and Feature Maps Generated-based Methods in CNNs
The MonoCNN improves the robustness of the CNN model via data-agnostic NLFs for filters but still with a complete convolution operation.
GhostNet [18]. The GhostNet introduces partial convolution filters ${\bm{W}}_{\mathrm{seed}}$ and generated the seed feature maps as follows:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{seed}}={\bm{W}}_{\mathrm{seed}}*{\bm{x}}\end{array},$ (5)
where ${\bm{y}}_{\mathrm{seed}}\in\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times c_{\mathrm{seed}}}$. For the remaining feature maps, GhostNet introduces additional cheap linear transformation, i.e., the $d\times d$ depthwise convolution filters to process the ${\bm{y}}_{\mathrm{seed}}$ for generating features:
${\bm{y}}$ $=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]$ (6)
$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{DWConv}}}\limits_{d\times d}({\bm{y}}_{\mathrm{seed}})]\ .$
However, the cheap linear operation (i.e., DWConv op.) introduced by GhostNet causes a slow computation in training. At the same time, the parameters of DWConv are also trainable, so the robustness of GhostNet is completely restricted to the training data.
SineFM [19]. Similar to the GhostNet, some filters of SineFM are designated as seed filters for a convolution layer. First, seed feature maps equal to the number of seed filters are generated based on the seed filters and input data. Then, feature maps with $t$-times the number of $c_{\mathrm{seed}}$ are generated based on the seed feature maps and the NLFs. Finally, using $1\times 1$ depthwise convolution for channel alignment.
Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input feature maps ${\bm{x}}$, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ of this layer by the Eq. 5. Then, $t$-data-agnostic NLFs $\bm{{\Psi}}_{\mathrm{sinefm}}={\{\Psi_{i}\|i=1,\dots,t\}}$ process the seed feature maps to obtain more features:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{exp}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{exp}_{1}};\cdots;{\bm{y}}_{\mathrm{exp}_{t}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[{\bm{y}}_{\mathrm{exp}_{i}}]\ ,\\ {\bm{y}}_{\mathrm{exp}_{i}}=\mathrm{\bf{Norm}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))\ ,\end{array}$ (7)
where ${\bm{y}}_{\mathrm{exp}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\!\times\!c_{\mathrm{seed}}}$, ${\bm{y}}_{\mathrm{exp}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\times\!t\cdot c_{\mathrm{seed}}}$, and $i\!=\!1,\!\cdots,\!t$. The $t$ is the times of expanding seed feature maps. The generated feature maps expanding by $t$ times of seed feature maps will usually exceed the number of feature maps required for this layer. Each expanding seed feature maps ${\bm{y}}_{\mathrm{exp}_{i}}$ also process a specific $\mathrm{\bf{Norm}}_{i}$. All expanding seed feature maps are concatenated and then process the $1\times 1$ depthwise convolution filters to align the dimensions and obtain the generated feature maps:
@@
$\begin{array}[]{l}{\bm{y}}_{\mathrm{gen}}=\mathop{\mathrm{\bf{DWConv}}}\limits_{1\times 1}({\bm{y}}_{\mathrm{exp}})\ .\end{array}$ (8)
@@
$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{DWConv}}}\limits_{1\times 1}(\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[\mathrm{\bf{Norm}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))])]\ .$
@@ 3.3 Design of GroupNL
3.3.1 Grouped Nonlinear Transformation Algorithm
The aforementioned GhostNet and SineFM methods introduce too much computation in feature maps generation by DWConv and Norm operations which result in more computational graph nodes and slow down training. At the same time, the data-agnostic nonlinear transformation of SineFM is only $t$-different types.
Therefore, resource efficiency and robustness are needed to be further improved. The key challenge is how to increase the diversity of data-agnostic nonlinear transformations to improve the robustness of the model without increasing the resource computation of running the model.
To this end, we propose the Grouped NonLinear transformation feature map generation (GroupNL) method, which is a robust and low-resource CNN model design method. Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input data/feature maps ${\bm{x}}$, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ of this layer as same as the Eq. 5.
Besides, we designate the seed filters ${\bm{W}}_{\mathrm{seed}}$ as the standard convolution in basic structures like ResNet-18/-34 and VGG11, and sparse in bottleneck structures, e.g., ResNet-101. The sparse filter means that the number of groups of the convolution operation of seed filters is the Greatest Common Divisor (GCD) of $c_{\mathrm{in}}$ and $c_{\mathrm{seed}}$, i.e., $\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. Introducing the sparse seed convolution in the larger bottleneck structure can further improve the resource efficiency of the larger CNN models. Utilizing the standard seed convolution in basic structure is beneficial for maintaining the performance of the smaller CNN models.
Splitting and Grouped Algorithm. We split the seed feature maps into $g$ groups in $c_{\mathrm{seed}}$ dimension instead of expanding, as follows:
@@ 3.3.1 Grouped Nonlinear Transformation Algorithm
where $g$ is a positive integer, and the modulo result is $\mathrm{Mod}(c_{\mathrm{seed}},g)\!=\!0$, and ${\bm{y}}_{\mathrm{seed}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times\frac{c_{\mathrm{seed}}}{g}},i=1,\cdots,g$. For the remaining dimension $c_{\mathrm{gen}}=c_{\mathrm{out}}-c_{\mathrm{seed}}$, we want to represent it by $c_{\mathrm{seed}}$ without expanding the dimension.
We denote a multiplier $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=(r\cdot c_{\mathrm{out}}/c_{\mathrm{in}}-1)$, so each group of seed feature maps should be copied $\gamma$ times for filling the remaining dimension. Thus, the total number of copied groups of feature maps is $\gamma\cdot g$.
Grouped Diverse Nonlinear Transformation. By splitting and copying, we have $\gamma\cdot g$ groups of identical seed feature maps, and no additional dimensions are introduced. We can introduce $\gamma\cdot g$ data-agnostic different nonlinear transformations, i.e., diverse NLFs $\bm{{\Psi}}_{\mathrm{groupnl}}={\{\Psi_{i,j}\|i=1,\dots,g;j=1,\dots,\gamma\}}$ to these identical groups. Subsequently, for each copied group of seed feature maps, NLF with different hyperparameter settings is used to generate the corresponding feature maps, as follows:
@@
As shown in Fig. 2, we can describe the operations of grouped diverse nonlinear transformation in detail with an example of $c_{\mathrm{in}}=8,c_{\mathrm{out}}=16,c_{\mathrm{seed}}=4,g=2$. In this case, the number of generated features $c_{\mathrm{gen}}=16-4=12$, the multiplier $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=12/4=3$, and introducing $|\bm{{\Psi}}_{\mathrm{groupnl}}|=\gamma\cdot g=6$ different NLFs for feature maps generation.
Finally, generated feature maps are concatenated with the seed feature maps to produce the output of the GroupNL layer, as follows:
@@
The PyTorch-like pseudocode of the GroupNL-based convolutional layer is demonstrated in Algorithm 1.
As described in Eq. 12 and Algorithm 1, the generating procedure of GroupNL is all in-placing (e.g., splitting $\tt{torch.split}$, concatenation $\tt{torch.cat}$) and copying (e.g., $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$) operators, and data-agnostic and random nonlinear transformation operations, without any extra $\tt{torch.nn.Module}$ (i.e., DWConv and Norm) compared with the GhostNet in Eq. 6 and SineFM in Eq. 9.
@@
class GroupNLConv2d(nn.Conv2d):
@@
# seed filters for generating feasures
self.conv = nn.Conv2d(C_i, C_s, \
@@
self.hy=nn.Parameter(torch.rand(gamma*g))
@@
Algorithm 1 GroupNL-based Convolutional Layer: PyTorch-like Pseudocode
3.3.2 Diversities of Data-Agnostic Nonlinear Transformation Function and Feature Maps Generation
@@
The number of NLFs in GroupNL is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g$, which is usually larger than the fixed number of $t$ NLFs in SineFM $|\Psi_{\mathrm{sinefm}}|=t$, which act on all seed feature maps for expanding, limiting the diversity of features.
@@
For the diversities of data-agnostic NLFs, assuming there are $64$ seed feature maps, and the number of output channels is $c_{\mathrm{out}}=256$, we can use different NLFs to generate the remaining $192$ feature maps. Assuming $t=5$, SineFM needs to generate $t\cdot c_{\mathrm{seed}}=5\cdot 64=320$ expanding features and utilizes a DWConv to align them into $192$ features. In this case, the number of NLFs is $5$. For GroupNL, the $\gamma=c_{\mathrm{gen}}/c_{\mathrm{seed}}=192/64=3$. For $g=2$, the number of different NLFs is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 2=6$; for $g=4$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 4=12$; for $g=8$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 8=24$; and for the fully grouped $g=64$ is $|\Psi_{\mathrm{groupnl}}|=\gamma\cdot g=3\cdot 64=192\gg|\Psi_{\mathrm{sinefm}}|=5$. Compared with the SineFM, the diversity of NLFs is significantly improved without increasing the number of feature maps in GroupNL.
@@
For the diversities of generated feature maps, we can visualize a given input image processing with 24-different NLFs as an example, as shown in Fig. 3. The richness of the generated feature maps is greatly improved while maintaining the similarity between different features.

### v2

@@ 3.1 Preliminaries
where * is the Conv operation. The MonoCNN improves the robustness of the CNN model via data-agnostic NLFs for filters, but still with a complete Conv operation.
GhostNet [11]. The Ghost Conv introduces $k\times k$ seed Conv to generate seed feature maps:
@@
$\begin{array}[]{l}{\bm{y}}_{\mathrm{seed}}={\bm{W}}_{\mathrm{seed}}*{\bm{x}}=\mathop{\mathrm{\bf{Conv}}}\limits_{\ k\times k}({\bm{x}})\end{array},$ (4)
where ${\bm{y}}_{\mathrm{seed}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times c_{\mathrm{seed}}}$. For the remaining feature maps, GhostNet introduces additional cheap linear transformation, i.e., the $d\!\times\!d$ cheap Conv to process the ${\bm{y}}_{\mathrm{seed}}$ for generating features:
${\bm{y}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};{\bm{y}}_{\mathrm{gen}}]=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{Conv}}}\limits_{d\times d}({\bm{y}}_{\mathrm{seed}})]\ .$ (5)
However, the cheap Conv introduced by GhostNet slows down training and inference. Also, the parameters of cheap Conv are trainable, so GhostNet’s robustness is completely restricted to the training data.
SineFM [27]. For a SineFM Conv, similar to the Ghost Conv, given seed filters ${\bm{W}}_{\mathrm{seed}}$ and input feature maps $\bm{x}$}, we can obtain seed feature maps ${\bm{y}}_{\mathrm{seed}}$ by the Eq. 4. Then, $t$-NLFs $\bm{{\Psi}}_{\mathrm{sinefm}}\!=\!{\{\Psi_{i}\|i\!=\!1,\!\dots\!,t\}}$ process the seed feature maps to obtain expanded feature maps ${\bm{y}}_{\mathrm{exp}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\!\times\!t\cdot c_{\mathrm{seed}}}$ with more features:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{exp}}=\mathrm{Concat}[{\bm{y}}_{\mathrm{exp}_{1}};\cdots;{\bm{y}}_{\mathrm{exp}_{t}}]=\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[{\bm{y}}_{\mathrm{exp}_{i}}]\ ,\\ {\bm{y}}_{\mathrm{exp}_{i}}=\mathrm{\bf{BN}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))\ ,\end{array}$ (6)
where ${\bm{y}}_{\mathrm{exp}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\!\times\!w^{\!{}^{\prime}\!}\!\times\!c_{\mathrm{seed}}}$, and $i\!=\!1,\!\cdots\!,\!t$. The $t$ is the time of expanding seed feature maps. The expanded feature maps will exceed the number of feature maps required for this layer. Each expanded seed feature maps ${\bm{y}}_{\mathrm{exp}_{i}}$ also process a specific $\mathrm{\bf{BN}}_{i}$. Then, expanded feature maps are processed by $1\!\times\!1$ cheap Conv to align the $c_{\mathrm{gen}}$ dimensions:
$\begin{array}[]{l}{\bm{y}}_{\mathrm{gen}}=\mathop{\mathrm{\bf{Conv}}}\limits_{1\times 1}({\bm{y}}_{\mathrm{exp}})\ .\end{array}$ (7)
@@
$=\mathrm{Concat}[{\bm{y}}_{\mathrm{seed}};\mathop{\mathrm{\bf{Conv}}}\limits_{1\times 1}(\mathop{\mathrm{Concat}}\limits_{i=1,\cdots,t}[\mathrm{\bf{BN}}_{i}(\Psi_{i}({\bm{y}}_{\mathrm{seed}}))])]\ .$
Figure 2: The example of GroupNL Conv with $c_{\mathrm{in}}\!=\!8,c_{\mathrm{out}}\!=\!16$, and the number of seed features is $c_{\mathrm{seed}}\!=\!4$, the groups for splitting is g=2g\!=\!2. In this case, the GroupNL can introduce $\gamma\cdot g\!=\!6$ different NLFs for the feature maps generation. Compared with the vanilla Conv, which needs 16 filters to compute all features, the GroupNL Conv only requires 4 filters to compute the seed features and generates the remaining features by nonlinearities and tensor manipulation operators.
@@ 3.2 Design of GroupNL
3.2.1 Grouped Nonlinear Transformation
The aforementioned GhostNet and SineFM methods introduce too much computation in feature maps generation by cheap Conv and BN operations, which result in slow training/inference speed. At the same time, NLFs in SineFM are only $t$-different types.
Therefore, resource efficiency and robustness need to be further improved. The key challenge is how to increase the diversity of data-agnostic NLFs to improve the robustness of the model without increasing computations.
To this end, we propose the Grouped NonLinear transformation (GroupNL) method. Formally, given the seed filters ${\bm{W}}_{\mathrm{seed}}$ and input feature maps $\bm{x}$}, we can obtain some seed feature maps ${\bm{y}}_{\mathrm{seed}}$ with the Eq. 4.
Splitting and Grouped Algorithm. We split seed feature maps into $g$ groups in $c_{\mathrm{seed}}$ dimension instead of expanding:
@@ 3.2.1 Grouped Nonlinear Transformation
where $g$ is a positive integer, where its modulo result is $\mathrm{Mod}(c_{\mathrm{seed}},g)\!=\!0$, and ${\bm{y}}_{\mathrm{seed}_{i}}\!\in\!\mathbb{R}^{h^{\!{}^{\prime}\!}\times{w^{\!{}^{\prime}\!}}\times\frac{c_{\mathrm{seed}}}{g}},i=1,\!\cdots\!,\!g$. For the remaining dimension $c_{\mathrm{gen}}\!=\!c_{\mathrm{out}}\!-\!c_{\mathrm{seed}}$, we want to represent it by $c_{\mathrm{seed}}$ without expanding the dimension.
We denote a multiplier $\gamma\!=\!\frac{c_{\mathrm{gen}}}{c_{\mathrm{seed}}}\!=\!r\cdot\frac{c_{\mathrm{out}}}{c_{\mathrm{in}}}-1$, so each group of seed feature maps should be copied $\gamma$ times for filling the remaining dimension. Thus, the total number of copied groups of feature maps is $\gamma\cdot g$.
Grouped Diverse Nonlinear Transformation. By splitting and copying, we have $\gamma\cdot g$ groups of identical seed feature maps, and no additional dimensions are introduced. We can introduce $\gamma\cdot g$ data-agnostic different nonlinear transformations, i.e., diverse NLFs $\bm{{\Psi}}_{\mathrm{gnl}}\!=\!{\{\Psi_{i,j}\|i\!=\!1,\dots,g;j\!=\!1,\dots,\gamma\}}$ to these identical groups. Subsequently, for each copied group of seed feature maps, NLFs with different hyperparameter settings are used to generate corresponding feature maps:
@@
As shown in Fig. 2, we can describe the operations of grouped diverse nonlinear transformation in detail with an example of $c_{\mathrm{in}}\!=\!8,c_{\mathrm{out}}\!=\!16,c_{\mathrm{seed}}\!=\!4,g\!=\!2$. In this case, the number of generated features $c_{\mathrm{gen}}\!=\!16\!-\!4\!=\!12$, the multiplier $\gamma\!=\!c_{\mathrm{gen}}/c_{\mathrm{seed}}\!=\!12/4\!=\!3$, and introducing $|\bm{{\Psi}}_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!6$ different NLFs for feature maps generation.
GroupNL Conv. Finally, generated feature maps are concatenated with the seed feature maps to produce the output of the standard GroupNL Conv layer, as follows:
@@
Sparse GroupNL Conv. In addition to the standard GroupNL Conv, we introduce a more lightweight module, termed Sparse GroupNL Conv. In this design, the seed filters ${\bm{W}}_{\mathrm{seed}}$ are grouped for thinning, where the number of groups $\xi$ in seed Conv is set to the Greatest Common Divisor (GCD) between $c_{\mathrm{in}}$ and $c_{\mathrm{seed}}$, i.e., $\xi\!=\!\mathrm{Gcd}(c_{\mathrm{in}},c_{\mathrm{seed}})$. The number of parameters, FLOPs, and energy consumption in Sparse GroupNL Conv are close to or slightly lower than depthwise Conv, while offering higher computational efficiency, as profiling in Section 4.4.3.
Therefore, Sparse GroupNL Conv can be used to replace depthwise Conv in lightweight models, e.g., MobileNet, EfficientNet, to improve the robustness.
Meanwhile, we observe that Sparse GroupNL Conv preserves performance in large-scale CNNs with bottleneck structures (e.g., ResNet-101), whereas the standard GroupNL Conv is still required for compact models based on basic blocks (e.g., ResNet-18/34).
The PyTorch-like pseudocode of the GroupNL Conv layer is demonstrated in Algorithm 1. As shown, its generating procedures are all tensor manipulation operators, e.g., splitting $\tt{torch.split}$, concatenation $\tt{torch.cat}$, copying $\tt{torch.repeat}$ and $\tt{torch.repeat\_interleave}$, and data-agnostic and random nonlinear transformation operations, without any extra $\tt{torch.nn.Module}$ (Conv and BN) compared with GhostNet in Eq. 5 and SineFM in Eq. 8.
@@
class GroupNLConv2d($n$.Conv2d):
@@
# seed filters for generating features
self.conv = $n$.Conv2d(C_i, C_s, \
@@
self.hy=$n$.Parameter(torch.rand(gamma*g))
@@
Algorithm 1 GroupNL Conv: PyTorch Pseudocode
3.2.2 Diversities of Nonlinearities and Feature Maps
@@
The number of NLFs in GroupNL is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g$, which is usually larger than the fixed number of $t$ NLFs in SineFM $|\Psi_{\mathrm{sinefm}}|\!=\!t$, which act on all seed feature maps to expand or limit feature diversity.
@@
Diversities of Data-agnostic NLFs. Assuming there are $64$ seed feature maps, and the number of output channels is $c_{\mathrm{out}}\!=\!256$, we can use different NLFs to generate the remaining $192$ feature maps.
For SineFM, assuming t=5t\!=\!5, it needs to generate $t\cdot c_{\mathrm{seed}}\!=\!5\cdot 64\!=\!320$ expanding features and utilizes a cheap Conv to align them into $192$ features. So, the number of NLFs is $5$.
For GroupNL, the $\gamma\!=\!c_{\mathrm{gen}}/c_{\mathrm{seed}}\!=\!192/64\!=\!3$. For $g=2$, the number of different NLFs is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 2\!=\!6$; for g=4g\!=\!4 is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 4\!=\!12$; for g=8g\!=\!8 is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 8\!=\!24$; and for the fully grouped g=64g\!=\!64 is $|\Psi_{\mathrm{gnl}}|\!=\!\gamma\cdot g\!=\!3\cdot 64\!=\!192\!\gg\!|\Psi_{\mathrm{sinefm}}|\!=\!5$.
Compared with the SineFM, the diversity of NLFs is significantly improved without increasing the number of feature maps in GroupNL.
@@
Diversities of Generated Feature Maps. We can visualize a given input image processing with 24-different NLFs as an example, as shown in Fig. 3. The richness of the generated feature maps is greatly improved while maintaining the similarity between different features.